# Detection and Characterization of Coordinated Online Behavior: A Survey

### v1

@@ 5.2. Harmfulness
Harmfulness measures the extent to which the actions of the coordinated actors can potentially cause negative consequences. Therefore, the analysis of the actions can provide valuable information about the intent, or potential, for harm of the coordinated actors. The production or re-sharing of content (e.g., text, images, links) is the most information-rich type of action and, in fact, nearly all indicators of harmfulness in Table 8 are derived from the analysis of content. A first line of work involves the use of traditional NLP methods to analyze attitudes, concerns, emotions, stances, topics, hashtags, and words, as these provide useful insights into the intent and aims of coordinated actors [14, 9, 26, 94, 91, 89, 133, 143, 135, 138, 139, 108, 24, 99, 62, 113, 143, 28, 134, 122, 88, 82]. For instance, certain words, negative emotions, and stances can amplify social discord and polarize public opinion. The prevalence of these indicators is largely due to the abundance of readily available tools and their ease of application. However, these indicators are often quite generic and consequently lack substantial power and informativeness. Text similarity scores were also used to identify harmful campaigns that aim to create false consensus through repeated sharing of similar posts [138, 139, 89, 60]. Other straightforward indicators of harmfulness are those based on the presence of toxic (i.e., hateful or offensive) [70, 108] or propagandistic [56] content. A different line of work measured harmfulness in terms of the unreliability of the news shared by coordinated actors, which was measured via NewsGuard’s [^4] https://www.newsguardtech.com/ (accessed: 31/07/2024) reliability ratings and Media Bias/Fact Check [^5] https://mediabiasfactcheck.com/ (accessed: 31/07/2024) factual reporting and credibility ratings [106, 121, 8, 9]. Others measured unreliability in terms of suspended and blacklisted URLs [9, 47, 135]. Finally, the fraction of coordinated users that were suspended or banned by a social platform has been extensively used as an indicator of harmfulness [56, 68, 67, 91, 99, 109, 121, 53, 141, 27, 113, 143, 88]. This stands out as the only indicator of harmfulness based on user, rather than content, information. However, we note that harmfulness indicators based on content moderation lack predictive capabilities and can only be used retrospectively, since moderation actions typically occur some time after the account have engaged in the harmful activities [126].
@@ 5.3. Orchestration
Orchestration expresses the extent to which the coordination is well-organized, rather than spontaneous and emergent. Since orchestration is related to the level of organization of a group of coordinated actors, all orchestration indicators are based on the analysis of coordination networks, as shown in Table 8. As such, these indicators are typically computed after the application of a network science detection method. A large coordination score among a group of users was used as a proxy for orchestration, based on the idea that tightly synchronized groups are likely to be well-organized [18, 56, 70, 90, 91, 89, 99, 121, 19, 66, 88]. Centrality scores measure the importance of nodes within a network or community. Hence, nodes with high centrality wield significant influence over the dissemination of information. As such, the presence of nodes with high centrality scores in a coordinated community could indicate a structured hierarchy [92, 48, 52, 66, 20, 29, 101, 64, 89, 27, 133, 88]. Conversely, assortativity is a measure of homophily in a network. Measuring assortativity can help determine whether coordinated behavior is orchestrated or spontaneous by assessing the extent to which highly connected nodes tend to connect with other highly connected nodes [19, 24, 99, 88]. For example, high assortativity between nodes with large degree might indicate a centralized orchestrated structure. Instead, high assortativity between nodes with low degree might indicate decentralized orchestration, while low assortativity or disassortativity might indicate spontaneous behaviors. Likewise, low modularity, high density, or high clustering coefficient indicates a tendency for nodes to form tightly-knit groups, suggesting well-organized and potentially orchestrated behavior [64, 96, 99, 48, 52, 92, 66, 89, 19, 27, 88]. Finally, the time intervals between the actions of coordinated actors can provide information on whether the coordinated behavior is orchestrated or spontaneous because tightly synchronized actions suggest a higher level of premeditated organization, whereas more irregular intervals might indicate spontaneous, less structured coordination. The study of the internal structure and organization of coordinated communities is also relevant for investigating the diverse strategies of online manipulation. Organized influence campaigns often unfold in distinct ways: some aim to mobilize pre-existing organic communities, leveraging existing social ties and trust, while others seek to establish new communities controlled by a central entity [118]. Additionally, organic communities can emerge from grassroots efforts, where users are genuinely motivated by a common goal without external orchestration. These different strategies leave “network footprints” that can be captured by orchestration indicators [19].
@@ 5.5. Other general indicators
In addition to the indicators that provide information about the four defining dimensions of coordinated behavior, other general-purpose indicators have also been used. These come in handy to provide additional context on the coordinated behavior and may be particularly relevant depending on the application context, as in the case of the indicators of political polarization or bias that are used to characterize coordinated communities involved in online electoral debates [70, 99, 66, 121, 122]. Political bias was computed based on MBFC scores or by leveraging hashtag polarity. Analyses of the level of activity of coordinated actors on a platform [14, 24, 26, 63, 67, 109, 110, 131, 135, 139, 45, 106, 133, 119, 62, 60], of the engagement they obtain [133, 106], or of the content they produce, can provide additional contextual information. Regarding the latter, standard socio-linguistic analyses were used to draw insights into the narratives and stances of coordinated actors [14, 9, 26, 94, 91, 89, 133, 143, 135, 138, 139, 108, 24, 99, 62, 113, 143, 134, 122, 88]. Similarly, others considered the diversity of the news shared by the coordinated actors [15, 48] or assessed the presence of patterns in the timings of their actions [12, 18, 101, 60]. Finally, acknowledging that coordination is a nuanced and non-binary concept, some studies computed coordination scores, for example based on the edge weights of the coordination network [18, 56, 70, 90, 91, 89, 99, 121, 19, 66, 88]. Analyzing the degree of coordination can contribute to shedding light on the extent of collaboration and influence within the group, and on the effectiveness of their collective actions.
@@ 5.6. Compound indicators
The previous discussion involved indicators used individually. However, some works also investigated the combined use of multiple indicators to derive even deeper insights, possibly revealing complex interactions and enhancing the overall understanding of coordinated behavior. For example, authors of [56, 99, 19, 121, 88] measured trends and correlations between indicators of coordination, propaganda, moderation, cohesiveness, and automation, finding that certain coordinated communities exhibit markedly different behaviors. They also found that assessments of the nature of online coordination—such as determining if it is harmful or inauthentic—can be unreliable when based on a single indicator, highlighting the importance of combining multiple indicators for more accurate results [56]. Similarly, others considered the interplay between coordination, toxicity, and political bias [70]. Finally, Tardelli et al. [121] proposed a comprehensive approach to characterize coordinated communities by analyzing multiple indicators. They examined coordination, automation, suspensions, news unreliability, and political bias, visualizing each community’s scores on these dimensions with a radar chart. They observed that larger scores across these dimensions indicate suspicious behavior. Therefore, communities with wider radar charts were deemed more suspicious than those with smaller areas. The area of a radar chart thus serves as a compound indicator, derived from combining the underlying dimensions. The suspiciousness of a coordinated community can thus be quantitatively assessed by calculating the radar chart’s area, or any other measure that correlates positively with the underlying dimensions. This approach exemplifies how multiple individual indicators can be combined to derive more detailed insights. Considering that there are only a few instances of the use of compound indicators in literature, this represents a largely unexplored area of research.
@@ 5.7. Discussion
Table 8 shows that many and diverse indicators are used to measure harmfulness. These include indicators of hate speech, toxicity, and propaganda, as well as indicators based on platform moderation decisions. Conversely, inauthenticity has only been examined through the lens of automation, despite being the focus of many works. From a practical standpoint, understanding the authenticity of an account is inherently more challenging than assessing its harmfulness, as authenticity pertains to the nature of the actors rather than their actions. Many online platforms afford users a significant degree of anonymity, making it difficult to verify their true identities. In contrast, many actions performed online leave unforgeable traces, which can be more easily analyzed to determine harmful behavior. Despite the challenges however, future work should focus on developing additional indicators of authenticity. Among the untapped sources of information are the completeness and credibility of the user profile, the regularity of activity patterns, the duplication or repetitiveness of posted content, the lack of consistency in language or interests, the authenticity of posted multimedia content, the cross-platform consistency of profile and behavior, the similarity of profile and behavior to that of previously moderated accounts, and the verification status, which however should be evaluated differently depending on whether such status can be purchased on the analyzed platform.666link.gale.com/apps/doc/A746844400/AONE (accessed: 31/07/2024)
@@ 6. Open challenges and future research directions
Some coordinated campaigns unfold simultaneously on multiple platforms resulting in multiplatform coordinated behavior [140]. These campaigns exhibit overall similar behaviors, such as the use of campaign-specific hashtags, with minor variations between platforms. In contrast, cross-platform campaigns involve different platforms playing distinct roles, with coordinated behavior varying significantly across the involved platforms. For example, organizers of a targeted hate attack might use private groups on a messaging platform to plan their actions before moving to the target platform to flood it with hateful comments [60]. Detecting coordination on the planning platform might involve analyzing group joinings, while on the target platform it could involve examining synchronized commenting. The example highlights the challenges of multi- and cross-platform analyses, which require extensive data collection, careful selection of actions or machine learning features for each platform, and identification of the same groups of users across multiple platforms. While the latter can be mitigated via the analysis of content rather than user networks—as discussed in Section 4.1.5—the remaining outstanding challenges result in a scarcity of multi- and cross-platform analyses. Figure 8 shows that most studies analyzed coordinated behavior on Twitter/X, with only a few multiplatform analyses [94, 95, 32, 106, 8, 144, 108, 60], highlighting that this area still necessitates much investigation.
Temporal variability.
@@
Most studies on coordinated behavior incorporate time to some degree, as shown in Tables 37, and 8. However, the majority of such works only perform superficial temporal analyses. Figure 8 highlights that only a few works have deeply investigated the temporal dynamics of coordinated behavior, using methods like multiplex temporal networks [122] and temporal point processes [113, 143]. Thus, analyzing temporal dynamics remains a promising yet relatively unexplored research avenue. Temporal analyses offer numerous advantages, including examining (i) the temporal stability of coordinated groups, (ii) changes in membership, structure, and actions, and (iii) adaptation to countermeasures or other stimuli [49]. Further, dynamic analyses seem to yield more accurate results compared to static ones [122]. However, this approach faces several challenges, such as setting appropriate temporal parameters and addressing the computational demands resulting from the time-intensive nature of dynamic analyses.
Heterogeneity.
@@
The great degree of heterogeneity inherent of coordinated behavior does not only manifests through temporal variations, but also through the multitude of characteristics that the different instances of coordinated behaviors exhibit. Figure 8 shows the distribution of studies that addressed specific types of online coordination, based on the defining dimensions introduced in Section 2.5. As shown by the vertical bars, coordinated harmful inauthentic and orchestrated behavior is the type of coordination that was studied the most. As discussed in Section 2, this is due to both practical reasons related to the urgency of contrasting online manipulations, as well as to the large body of work that adopted Facebook’s initial Definition 2.2 of coordinated inauthentic behavior. Almost all the other remaining combinations involve harmful behavior and static analyses. Figure 8 also shows that authentic spontaneous and harmless coordinated behavior is completely unstudied. While harmful, inauthentic, and orchestrated behaviors are of particular practical relevance for content moderation purposes, the study of harmless, authentic, and spontaneous coordination should not be overlooked as it provides valuable insights into fundamental aspects of online human interactions. Therefore, the analysis of the less problematic forms of online coordination should be increased in the near future.
Multimodality.
@@
Online coordination can involve multiple content modalities, such as text, images, and videos. This diversity of content types adds complexity to detecting and analyzing coordinated behavior, as each modality has unique characteristics requiring different analytical techniques. For instance, text-based coordination might involve analyzing text similarities or hashtags use [133, 27, 135]. Image-based coordination could require image recognition tasks or computing image similarities [141, 46, 102], while video-based coordination demands techniques like video content analysis and transcript examination. As shown in Tables 26, and 7, however, only a handful of works investigated modalities other than text [84, 102, 96, 141, 116, 46]. Accounting for and integrating different modalities offers the potential to capture a more comprehensive picture of coordinated behavior, but also poses new challenges. These include developing features and methodologies to assess similarities based on multimedia content and addressing the computational demands of multimodal analyses. Nonetheless, given the rising trend of online platforms centered around multimedia content, multimodal analyses appear promising and well motivated.
Generative AI
@@
The rise of generative AI presents both challenges and opportunities. One potential challenge is the increased difficulty in detecting coordinated actors that use such techniques to mask their activities. AI is already capable of producing human-like text with given properties, and increasingly also images, audio, and video [97]. Moreover, it has already been used to simulate authentic online human behaviors in such a way as to evade detection by existing systems [54]. Progress in generative AI could thus make it harder to detect future instances of coordinated behavior or to assess the authenticity of coordinated actors. However, it is still unknown what the effect of these techniques will be on the landscape of online coordination. At the same time, generative AI (e.g., LLMs) can be used to simulate human behaviors within agent-based models [123], enabling the creation of realistic simulations of coordinated activities. In future, these models could be leveraged to train or evaluate detection methods, providing a controlled environment to probe methods’ capabilities at detecting different instances and types of coordinated behavior.
Scalability.
@@
Large-scale coordinated campaigns often involve tens of thousands of users and millions of interactions, creating a heavy computational burden. This issue is especially pronounced for methods that rely on computing pairwise similarities between users, such as the majority of network science methods. Consequently, some studies analyze only a tiny fraction of users—sometimes as little as 1% [99, 29]—which limits the scope and reliability of coordination detection results. Future efforts should focus on increasing the scalability of current detection methods to handle the vast amounts of data generated by large-scale campaigns. Possible solutions to this issue can include the use of approximation algorithms to obtain near-accurate results with reduced computational complexity, or sampling techniques to select representative data subsets, making computations more manageable. Additionally, online algorithms could be used to update results continuously as new data arrives rather than recomputing from scratch, and graph summarization techniques could be used to condense large networks while preserving key structural properties for faster analysis.
Data availability.
The scarcity of labeled authoritative data limits the development of new coordination detection methods. This constraint hampers the creation of supervised methods, leading to the predominance of unsupervised approaches that we documented in Section 4. Moreover, it complicates the fine-tuning and evaluation of all methods, now primarily assessed through post-hoc characterizations rather than via quantitative metrics. Unfortunately, the challenge of data availability is expected to intensify in the future, as access to platforms data continues to diminish in the post-API age [125, 61]. Potential solutions include using certain public data as ground truth or running simulations. Examples of the former approach are those works that built balanced ground-truths based on authoritative and publicly available repositories, [^7] https://transparency.x.com/en/reports/moderation-research.html (accessed: 31/07/2024) obtaining datasets that are suitable for training [35, 73] and evaluation [122] purposes. Existing datasets are however constrained to only one type of coordinated behavior on a single platform, which is insufficient for training detectors capable of generalizing to the many instances of online coordination, such as those sketched in Figure 3. This calls for further endeavors to create more comprehensive ground-truths. An alternative solution is to evaluate detection methods in simulated environments, as done in recent studies that employed agent-based models to mimic coordinated behavior [87, 104, 59]. In any case, ensuring access to reliable benchmarks—either real or synthetic—will remain a critical issue that demand significant attention in the coming years, particularly given the dynamic and complex nature of coordinated online behavior, which necessitate diverse and frequently updated data.
Subjectivity.
@@
The existence of many indicators of harmfulness—reported in Table 8—is advantageous from a practical standpoint. However, this multitude of indicators reflects great variability in how harmfulness has been practically defined. More broadly, the subjective nature of harmfulness—discussed in Sections 2.4.3 and 2.5.2—presents several challenges to its assessment. The first of such challenges is the variability with which different stakeholders define what constitutes harmful behavior, which hinder reaching a consensus on definitions and indicators [31]. Additionally, perceptions of harmfulness are influenced by cultural and social norms, which differ across individuals, regions, and communities. Consequently, what is seen as harmful by some might not be perceived the same way by others. Finally, assessing harmfulness involves evaluating both the intent behind the actions and their impact. While the impact can sometimes be measured—as with misinformation that leads to real-world violence [134] or that reduces vaccine uptake [69]—the intent is often hidden or ambiguous, making it difficult to measure accurately. Given that a certain degree of subjectivity in the assessment of online coordination is inevitable and bound to persist, future works must prioritize transparency in defining and implementing it, favoring nuanced analyses based on multiple indicators.
Attribution.
@@
Detecting coordination, especially in cases where malicious actors aim to remain hidden, is already challenging. Attributing these instances to the responsible entities is even more difficult. Attribution does not merely involve identifying the accounts that take part in the coordination, but most importantly uncovering the entities behind them—such as movements, organizations, or states—and understanding their goals. Several obstacles hinder this process, including the anonymity afforded by online platforms and the sophisticated techniques used to obfuscate identities, activities, and intentions. Potential solutions include strengthening the collaboration with platforms, which have access to more data than what is publicly visible or available via APIs or scraping. Additionally, advanced forensic techniques and multidisciplinary approaches that combine technical, social, and behavioral analysis are essential. Despite these solutions, the task will remain daunting and require continuous research and experimentation.
Multidisciplinarity.
@@
The study of coordinated behavior is intrinsically multidisciplinary due to the complex interplay of technical, social, and regulatory factors involved. This multidisciplinary nature is evident in the diversity of the studies on the subject, which employ a wide array of methods from multiple scientific communities. For instance, computer scientists and complex systems scholars develop computational methods to detect instances of coordinated behavior [121, 75, 139]. Social scientists study the emergent behaviors and interactions that lead to massive online coordination [26, 66], while political scientists analyze the impact of coordinated campaigns on real-world events, such as elections [99, 18, 141]. The involvement of media and communication experts in understanding the dissemination of (mis)information further highlights the multidisciplinary scope [116, 48]. In this context, multidisciplinarity is both a challenge and an opportunity. Challenges arise from the need to coordinate efforts among various scientific communities and to organize and keep track of the extensive body of work on the subject—a task for which this survey aims to provide a contribution. In spite of the challenges however, the comprehensive understanding of the phenomenon and the effective mitigation of its nefarious instances can only be achieved through the collaboration between diverse stakeholders, such as scholars, platform operators, policymakers, and civil society organizations.
Ethical dilemmas.

### v2

@@ 5.2. Harmfulness
Harmfulness measures the extent to which the actions of the coordinated actors can potentially cause negative consequences. Hence, the analysis of the actions can provide valuable information about the intent, or potential, for harm of the coordinated actors. The production or re-sharing of content (e.g., text, images, links) is the most information-rich type of action and, in fact, nearly all indicators of harmfulness in Table 9 are derived from the analysis of content. Traditional NLP methods are used to analyze attitudes, concerns, emotions, stances, topics, hashtags, and words, as these provide useful insights into the intent and aims of coordinated actors [15, 12, 26, 97, 94, 91, 140, 155, 142, 145, 146, 113, 24, 100, 58, 118, 155, 29, 90, 80, 13, 20, 122, 149, 136]. For instance, certain words, negative emotions, and stances can amplify social discord and polarize public opinion. The prevalence of these indicators is largely due to the abundance of readily available tools and their ease of application. However, these indicators are often quite generic and consequently lack substantial power and informativeness. Text similarity scores were also used to identify harmful campaigns that aim to create false consensus through repeated sharing of similar posts [145, 146, 91, 55]. Other straightforward indicators of harmfulness are those based on the presence of toxic (i.e., hateful or offensive) [66, 113] or propagandistic [51, 84] content. A different line of work measured harmfulness in terms of the unreliability of the news shared by coordinated actors, which was measured via NewsGuard’s [^7] https://www.newsguardtech.com/ (accessed: 31/07/2024) reliability ratings and Media Bias/Fact Check [^8] https://mediabiasfactcheck.com/ (accessed: 31/07/2024) factual reporting and credibility ratings [110, 128, 11, 12, 20, 122, 150]. Others measured unreliability in terms of suspended and blacklisted URLs [142, 43, 12, 20]. The fraction of coordinated users that were suspended or banned by a social platform has been extensively used as an indicator of harmfulness [51, 65, 64, 94, 100, 115, 128, 49, 152, 28, 118, 155, 90]. Finally, the use of AI generated content may be a signal of potentially harmful activity [20]. This stands out as the only indicator of harmfulness based on user, rather than content, information. However, we note that harmfulness indicators based on content moderation lack predictive capabilities and can only be used retrospectively, since moderation actions typically occur some time after the account have engaged in the harmful activities.
@@ 5.3. Orchestration
Orchestration expresses the extent to which the coordination is well-organized, rather than spontaneous and emergent. Since orchestration reflects the level of organization among coordinated actors, its indicators are derived from coordination networks (Table 9) and typically computed after network-based detection. High coordination scores are used as proxies for orchestration, under the assumption that tightly synchronized groups are more organized [17, 51, 66, 93, 94, 91, 100, 128, 18, 62, 90, 84]. Centrality scores measure node importance within a network or community. Highly central nodes have a greater influence on the dissemination of information, and their presence in coordinated communities may indicate a structured hierarchy [95, 44, 48, 62, 19, 30, 102, 60, 91, 28, 140, 90]. Conversely, assortativity is a measure of homophily in a network. Measuring assortativity can help determine whether coordinated behavior is orchestrated or spontaneous by assessing the extent to which highly connected nodes tend to connect with other highly connected nodes [18, 24, 100, 90, 149]. For example, high assortativity between nodes with large degree might indicate a centralized orchestrated structure. Instead, high assortativity between nodes with low degree might indicate decentralized orchestration, while low assortativity or disassortativity might indicate spontaneous behaviors. Likewise, low modularity, high density, or high clustering coefficient indicates a tendency for nodes to form tightly-knit groups, suggesting well-organized and potentially orchestrated behavior [60, 99, 100, 44, 48, 95, 62, 91, 18, 28, 90, 122]. Finally, the time intervals between the actions of coordinated actors can provide information on whether the coordinated behavior is orchestrated or spontaneous because tightly synchronized actions suggest a higher level of premeditated organization, whereas more irregular intervals might indicate spontaneous, less structured coordination. The study of the internal structure and organization of coordinated communities is also relevant for investigating the diverse strategies of online manipulation. Organized influence campaigns often unfold in distinct ways: some aim to mobilize pre-existing organic communities, leveraging existing social ties and trust, while others seek to establish new communities controlled by a central entity [125]. Additionally, organic communities can emerge from grassroots efforts, where users are genuinely motivated by a common goal without external orchestration. These different strategies leave “network footprints” that can be captured by orchestration indicators [18].
@@ 5.5. Other general indicators
In addition to the indicators that provide information about the four defining dimensions of coordinated behavior, other general-purpose indicators have also been used. These come in handy to provide additional context on the coordinated behavior and may be particularly relevant depending on the application context, as in the case of the indicators of political polarization or bias that are used to characterize coordinated communities involved in online electoral debates [66, 100, 62, 128, 129]. Political bias was computed based on MBFC scores or by leveraging hashtag polarity. Analyses of the level of activity of coordinated actors on a platform [15, 24, 26, 59, 64, 115, 116, 138, 142, 146, 41, 110, 140, 126, 58, 55], of the engagement they obtain [140, 110], or of the content they produce, can provide additional contextual information. Regarding the latter, standard socio-linguistic analyses were used to draw insights into the narratives and stances of coordinated actors [15, 12, 26, 97, 94, 91, 140, 155, 142, 145, 146, 113, 24, 100, 58, 118, 155, 141, 129, 90, 122, 61]. Similarly, others considered the diversity of the news shared by the coordinated actors [16, 44] or assessed the presence of patterns in the timings of their actions [14, 17, 102, 55]. Finally, acknowledging that coordination is a nuanced and non-binary concept, some studies computed coordination scores, for example based on the edge weights of the coordination network [17, 51, 66, 93, 94, 91, 100, 128, 18, 62, 90, 63].Analyzing the degree of coordination helps reveal the level of collaboration, influence, and effectiveness of collective actions within a group.
@@ 5.6. Compound indicators
While the previous discussion considered indicators individually, combining multiple indicators can provide deeper insights by capturing complex interactions. For instance, studies [51, 100, 18, 128, 90] analyze trends and correlations across dimensions such as coordination, propaganda, moderation, cohesiveness, and automation, showing that coordinated communities can exhibit distinct behaviors. They also highlight that relying on a single indicator may lead to unreliable assessments, emphasizing the importance of compound indicator approaches [51]. Similarly, others considered the interplay between coordination, toxicity, and political bias [66]. In [128], a comprehensive characterization of coordinated communities is proposed using multiple indicators, including coordination, automation, suspensions, news unreliability, and political bias, visualized via radar charts. Higher scores—and thus larger chart areas—indicate more suspicious behavior, making the radar area a compound indicator of suspiciousness. This approach highlights the potential of combining multiple indicators, an area that remains largely underexplored.
@@ 5.7. Discussion
Table 9 shows that many and diverse indicators are used to measure harmfulness. These include indicators of hate speech, toxicity, and propaganda, as well as indicators based on platform moderation decisions. Conversely, inauthenticity has only been examined through the lens of automation, despite being the focus of many works. From a practical standpoint, understanding the authenticity of an account is inherently more challenging than assessing its harmfulness, as authenticity pertains to the nature of the actors rather than their actions. Many online platforms afford users a significant degree of anonymity, making it difficult to verify their true identities. In contrast, many actions performed online leave unforgeable traces, which can be more easily analyzed to determine harmful behavior. Despite the challenges however, future work should focus on developing additional indicators of authenticity. Among the untapped sources of information are the completeness and credibility of the user profile, the regularity of activity patterns, the duplication or repetitiveness of posted content, the lack of consistency in language or interests, the authenticity of posted multimedia content, the cross-platform consistency of profile and behavior, the similarity of profile and behavior to that of previously moderated accounts, and the verification status, which however should be evaluated differently depending on whether such status can be purchased on the analyzed platform.999link.gale.com/apps/doc/A746844400/AONE (accessed: 31/07/2024)
@@ 6. Open challenges and future research directions
Multiplatform-ness. Some coordinated campaigns unfold simultaneously on multiple platforms, resulting in multiplatform coordinated behavior [147]. These campaigns exhibit overall similar behaviors, such as the use of campaign-specific hashtags, with minor variations between platforms. In contrast, cross-platform campaigns involve different platforms playing distinct roles, with coordinated behavior varying significantly across the involved platforms. For example, organizers of a targeted hate attack might use private groups on a messaging platform to plan their actions before moving to the target platform to flood it with hateful comments [55]. Detecting coordination on the planning platform might involve analyzing group joinings, while on the target platform it could involve examining synchronized commenting. The example highlights the challenges of multi- and cross-platform analyses, which require extensive data collection, careful selection of actions or machine learning features for each platform, and identification of the same groups of users across multiple platforms. While the latter can be mitigated via the analysis of content rather than user networks—as discussed in Section 4.1.5—the remaining outstanding challenges result in a scarcity of multi- and cross-platform analyses. Figure 10 shows that most studies analyzed coordinated behavior on Twitter/X, with only a few multiplatform analyses [97, 98, 32, 110, 11, 156, 113, 55, 81, 52], highlighting that this area still necessitates much investigation.
@@
Temporal variability. Most studies on coordinated behavior incorporate time to some degree, as shown in Tables 378, and 9. However, the majority of such works only perform superficial temporal analyses. Figure 10 highlights that only a few works have deeply investigated the temporal dynamics of coordinated behavior, using methods like multiplex temporal networks [129] and temporal point processes [118, 155]. Thus, analyzing temporal dynamics remains a promising yet relatively unexplored research avenue. Temporal analyses offer numerous advantages, including examining (i) the temporal stability of coordinated groups, (ii) changes in membership, structure, and actions, and (iii) adaptation to countermeasures or other stimuli [45]. Further, dynamic analyses seem to yield more accurate results compared to static ones [129]. However, this approach faces several challenges, such as setting appropriate temporal parameters and addressing the computational demands resulting from the time-intensive nature of dynamic analyses.
@@
Heterogeneity. The great degree of heterogeneity inherent of coordinated behavior does not only manifests through temporal variations, but also through the multitude of characteristics that the different instances of coordinated behaviors exhibit. Figure 10 shows the distribution of studies that addressed specific types of online coordination, based on the defining dimensions introduced in Section 2.5. As shown by the vertical bars, coordinated harmful inauthentic and orchestrated behavior is the type of coordination that was studied the most. As discussed in Section 2, this is due to both practical reasons related to the urgency of contrasting online manipulations, as well as to the large body of work that adopted Facebook’s initial Definition 2.2 of coordinated inauthentic behavior. Almost all the other remaining combinations involve harmful behavior and static analyses. Figure 10 also shows that authentic spontaneous and harmless coordinated behavior is completely unstudied. While harmful, inauthentic, and orchestrated behaviors are of particular practical relevance for content moderation purposes, the study of harmless, authentic, and spontaneous coordination should not be overlooked as it provides valuable insights into fundamental aspects of online human interactions. Therefore, the analysis of the less problematic forms of online coordination should be increased in the near future.
@@
Multimodality. Online coordination can involve multiple content modalities, such as text, images, and videos. This diversity of content types adds complexity to detecting and analyzing coordinated behavior, as each modality has unique characteristics requiring different analytical techniques. For instance, text-based coordination might involve analyzing text similarities or hashtags use [140, 28, 142]. Image-based coordination could require image recognition tasks or computing image similarities [152, 42, 103], while video-based coordination demands techniques like video content analysis and transcript examination. As shown in Tables 2678 however, only a handful of works investigated modalities other than text [83, 103, 99, 152, 121, 42]. Accounting for and integrating different modalities offers the potential to capture a more comprehensive picture of coordinated behavior, but also poses new challenges. These include developing features and methodologies to assess similarities based on multimedia content and addressing the computational demands of multimodal analyses. Nonetheless, given the rising trend of online platforms centered around multimedia content, multimodal analyses appear promising and well motivated.
@@
Generative AI The rise of generative AI presents both challenges and opportunities. One potential challenge is the increased difficulty in detecting coordinated actors that use such techniques to mask their activities. AI is already capable of producing human-like text with given properties, and increasingly also images, audio, and video [20]. Moreover, it has already been used to simulate authentic online human behaviors in such a way as to evade detection by existing systems [96]. Progress in generative AI could thus make it harder to detect future instances of coordinated behavior or to assess the authenticity of coordinated actors. However, it is still unknown what the effect of these techniques will be on the landscape of online coordination. At the same time, as discussed in Section 4.4, recent work has begun to explore LLM-driven agents to simulate coordinated behavior in controlled environments [101]. Looking ahead, these approaches could enable more realistic and scalable simulations, supporting the development and stress-testing of detection methods. However, the use of LLMs in this context is still in its early stages, and a systematic understanding of their capabilities and limitations for modeling coordinated behavior is currently lacking.
@@
Scalability. Large-scale coordinated campaigns can involve tens of thousands of users and millions of interactions, posing significant computational challenges, especially for methods based on pairwise user similarities. As a result, some studies analyze only a small fraction of data—sometimes as little as 1% [100, 30]—limiting reliability. Future efforts should focus on increasing the scalability of current detection methods to handle the vast amounts of data generated by large-scale campaigns. Possible solutions to this issue can include the use of approximation algorithms to obtain near-accurate results with reduced computational complexity, or sampling techniques to select representative data subsets. Additionally, online algorithms can update results incrementally as new data arrives, while graph summarization techniques can condense large networks while preserving key structural properties.
Validation and formalization. As discussed in Section 4.4, current validation strategies only partially alleviate the lack of labeled data. The challenge of data availability is expected to intensify as access to platform data continues to decline in the post-API era [132]. Although some works leverage publicly available repositories to construct ground-truth datasets [70, 81, 18], these resources remain limited in scope, typically focusing on a single platform or a specific type of coordinated behavior. This restricts the ability of models to generalize across the diverse manifestations of coordination, such as those illustrated in Figure 5. At the same time, simulation-based solutions are still underexplored in the coordinated behavior literature [86, 53, 101]. As discussed in Section 2.3, the field lacks a unified and statistically grounded definition of coordinated behavior, which limits principled validation. A key challenge is to translate existing operational definitions into formal, testable frameworks, for instance through appropriate null models capturing baseline behavior. Such models would enable more robust and comparable assessments across datasets and domains. In this direction, theory-driven approaches offer a promising path, drawing from socio-psychological models of collective action [137] and the statistical physics of complex networks [2] to develop more general and interpretable frameworks.
@@
Subjectivity. The existence of many indicators of harmfulness—reported in Table 9—is advantageous from a practical standpoint. However, this multitude of indicators reflects great variability in how harmfulness has been practically defined. More broadly, the subjective nature of harmfulness—discussed in Sections 2.4.3 and 2.5.2—presents several challenges to its assessment. The first of such challenges is the variability with which different stakeholders define what constitutes harmful behavior, which hinder reaching a consensus on definitions and indicators [31]. Additionally, perceptions of harmfulness are influenced by cultural and social norms, which differ across individuals, regions, and communities. Consequently, what is seen as harmful by some might not be perceived the same way by others. Finally, assessing harmfulness involves evaluating both the intent behind the actions and their impact. While the impact can sometimes be measured—as with misinformation that leads to real-world violence [141] or that reduces vaccine uptake [122]—the intent is often hidden or ambiguous, making it difficult to measure accurately. Given that a certain degree of subjectivity in the assessment of online coordination is inevitable and bound to persist, future works must prioritize transparency in defining and implementing it, favoring nuanced analyses based on multiple indicators.
@@
Attribution. Detecting coordination, especially in cases where malicious actors aim to remain hidden, is already challenging. Attributing these instances to the responsible entities is even more difficult. Attribution does not merely involve identifying the accounts that take part in the coordination, but most importantly uncovering the entities behind them—such as movements, organizations, or states—and understanding their goals. Several obstacles hinder this process, including the anonymity afforded by online platforms and the sophisticated techniques used to obfuscate identities, activities, and intentions. Potential solutions include strengthening the collaboration with platforms, which have access to more data than what is publicly visible or available via APIs or scraping. Additionally, advanced forensic techniques and multidisciplinary approaches that combine technical, social, and behavioral analysis are essential. Despite these solutions, the task will remain daunting and require continuous research and experimentation.
@@
Multidisciplinarity. The study of coordinated behavior is intrinsically multidisciplinary due to the complex interplay of technical, social, and regulatory factors involved. This multidisciplinary nature is evident in the diversity of the studies on the subject, which employ a wide array of methods from multiple scientific communities. For instance, computer scientists and complex systems scholars develop computational methods to detect instances of coordinated behavior [128, 73, 146]. Social scientists study the emergent behaviors and interactions that lead to massive online coordination [26, 62], while political scientists analyze the impact of coordinated campaigns on real-world events, such as elections [100, 17, 152]. The involvement of media and communication experts in understanding the dissemination of (mis)information further highlights the multidisciplinary scope [121, 44]. In this context, multidisciplinarity is both a challenge and an opportunity. Challenges arise from the need to coordinate efforts among various scientific communities and to organize and keep track of the extensive body of work on the subject—a task for which this survey aims to provide a contribution. In spite of the challenges however, the comprehensive understanding of the phenomenon and the effective mitigation of its nefarious instances can only be achieved through the collaboration between diverse stakeholders, such as scholars, platform operators, policymakers, and civil society organizations.